# Online Price Competition under Generalized Linear Demands

### v3

@@ Abstract
We study sequential price competition among $N$ sellers, each influenced by the pricing decisions of their rivals. Specifically, the demand function for each seller $i$ follows the single index model $\lambda_{i}(\mathbf{p})=\mu_{i}(\langle\boldsymbol{\theta}_{i,0},\mathbf{p}\rangle)$, with known increasing link $\mu_{i}$ and unknown parameter $\boldsymbol{\theta}_{i,0}$, where the vector $\mathbf{p}$ denotes the vector of prices offered by all the sellers simultaneously at a given instant. Each seller observes only their own realized demand – unobservable to competitors – and the prices set by rivals. Our framework generalizes existing approaches that focus solely on linear demand models. We propose a novel decentralized policy, PML-GLUCB, that combines penalized MLE with an upper-confidence pricing rule, removing the need for coordinated exploration phases across sellers – which is integral to previous linear models – and accommodating both binary and real-valued demand observations. Relative to a dynamic benchmark policy, each seller achieves $O(N^{2}\sqrt{T}\log(T))$ regret, which essentially matches the optimal rate known in the linear setting. A significant technical contribution of our work is the development of a variant of the elliptical potential lemma – typically applied in single-agent systems – adapted to our competitive multi-agent environment.
@@ 1 Introduction
Pricing is the strategic engine of competitive markets: firms revise prices not only to track shifting demand, but to anticipate and counter rivals’ moves. Algorithmic pricing amplifies this feedback, placing learning at the center of price formation. Yet much of the learning-to-price literature still leans on monopolistic or weakly interactive assumptions (fan2024policy; javanmard2019dynamic; golrezaei2019dynamic; bracale2025dynamic), effectively treating demand as if competitors were irrelevant. Recent works have designed policies in multiple seller interaction scenarios, where each seller is influenced by their competitors’ strategies. These works propose sequential pricing mechanisms that achieve sublinear regret [^1] In this paper, we use the terms sublinear regret and no regret interchangeably to indicate that the cumulative regret grows sub-linearly with time, that is, $\mathrm{Reg}(T)=o(T)$ as $T\to\infty$. and converge to a Nash equilibrium (NE) (li2024lego; bracale2025revenue). Yet two issues remain for practice: the demand is often forced to be linear (or narrowly parameterized) (li2024lego; goyal2023learning), and many methods separate “explore” from “optimize,” with phase lengths that depend on unknown constants (li2024lego; bracale2025revenue) and i.i.d. pricing policy in the exploration phase.
“A natural question arises: can a seller update individual parameter estimators in each period and subsequently compute an optimal price to offer based on their current estimates? […] Literature in both statistics and operations research has demonstrated that such a myopic pricing policy does not necessarily lead to convergence to the Nash equilibrium or individual optimal prices.” (li2024lego)
In this paper, we address this question within a more flexible and realistic model of competition. Specifically, we adopt a generalized linear (single-index) demand framework to capture the broader relevance of nonlinear demand patterns observed in real-world markets (gallego2006price; wan2022nonlinear; wan2023conditions). Building on this model, we propose a myopic pricing-while-estimating policy that operates without a dedicated i.i.d. pricing design exploration phase. Despite the absence of such a phase and, even if the resulting price trajectory does not converge to a Nash equilibrium (NE), we prove that the individual regret still scales as $\widetilde{O}(\sqrt{T})$ against a dynamic benchmark, where $\widetilde{O}$ suppresses logarithmic factors.
It is important to emphasize that in this work, we do assume that a NE does exist; however, the convergence of our algorithm to this equilibrium is not a prerequisite for achieving the sublinear regret guarantee established in this work.
@@ 1.1 Related Literature
Over the past two decades, there has been substantial progress in demand learning under competition. Early work by kirman1975learning analyzed symmetric two-seller competition with linear demand. Asymmetry and noise were incorporated by cooper2015learning, who let each seller learn independently (effectively as a monopolist) and established equilibrium under known price sensitivities. In contrast, li2024lego studied dynamic pricing with unknown price sensitivities under a linear demand model and achieved the optimal $\widetilde{O}(\sqrt{T})$ regret rate. In their setup, firm $i$’s expected demand is
@@
where $N]=\{1,2,\dots,N\}$ indexes the sellers and the parameters $(\alpha_{i,0},\beta_{i,0},\gamma_{ij,0})$ are unknown with $\beta_{i,0}>0$. The condition $\beta_{i,0}>0$ captures the standard economic assumption that the demand of seller $i$ decreases as their own price $p_{i}$ increases. However, nonlinear demand models are often more realistic and have been explored in several applications (gallego2006price; wan2022nonlinear; wan2023conditions). For instance, gallego2006price considered the mean demand function $\lambda_{i}(\mathbf{p})=a_{i}\left(p_{i}\right)/(\sum_{j}a_{j}\left(p_{j}\right)+\kappa)$, for some known increasing functions $a_{i}$ and real value $\kappa\in(0,1]$. They establish the existence of a NE, but they do not propose an online learning procedure.
In contrast to prior studies (li2024lego; kachani2007modeling; gallego2006price), we analyze a more general monotone single-index expected-demand model:
@@
for each $i\in[N]$, where $\mathbf{p}_{-i}=(p_{j})_{j\in[N]\setminus\{i\}}$, $\mathbf{p}=(p_{j})_{j\in[N]}$ and $\boldsymbol{\theta}_{i,0}$ is a vector of dimension $N$ with $i$-th entry equals to $\beta_{i,0}$ (where $\beta_{i,0}>0$) and the rest of the entries are the values of $\boldsymbol{\gamma}_{i,0}$ ordered. The link function $\mu_{i}$ is assumed to be increasing, capturing the same economic intuition that demand decreases with the seller’s own price. Here $\mu_{i}$ is known to seller $i$ (but unknown to competitors) and the parameter $\boldsymbol{\theta}_{i,0}$ is unknown (to everyone). This model generalizes the linear demand model by li2024lego; specifically, $\mu_{i}(u)=u+\alpha_{i,0}$ recovers their formulation.
@@
A key modeling assumption in MARL algorithms is full or partial observability of other agents’ rewards or payoffs (hu2003nash; zhang2021multi; yang2020overview). For instance, the canonical formulation of hu2003nash assumes that each agent can observe the immediate reward of all players and use this to update joint policy. Similarly, zhang2021multi and yang2020overview rely on either joint reward observability or access to sufficient statistics (e.g., payoff signals or utility vectors) that allow agents to infer the equilibrium structure. This assumption is critical: observing other agents’ realized rewards enables an agent to estimate the full payoff landscape and to learn best-response dynamics with respect to the joint reward function.
Our setting fundamentally differs from this MARL paradigm in both information structure and observability. In the sequential pricing game we study, each seller observes only their own reward (or demand value) and rivals’ posted prices – never competitors’ demands or revenues. Because demand functions are private and unobservable, a seller cannot recover rivals’ payoff functions or utility gradients, so classical MARL techniques are not directly applicable. Nevertheless, we design a policy whose dynamics need not converge to a Nash equilibrium, yet whose individual regret satisfies $\widetilde{O}(\sqrt{T})$. This makes our approach better aligned with real-world competitive pricing, where revenue data is proprietary and competitors’ demand is hidden.
@@
One may view $\mathbf{p}_{-i}^{(t)}$ at any instant $t$ as the “context” for seller $i$, linking our problem to contextual bandits (chu2011contextual; agrawal2013thompson; chakraborty2023thompson; kim2023double) and to the nonlinear Cournot structure (cournot1838recherches; saloner1987cournot). However, unlike standard contextual bandits – where the context is observed before the action is taken – in our market the sellers post prices simultaneously and observe $\mathbf{p}_{-i}^{(t)}$ only after choosing $p_{i}^{(t)}$. This endogeneity/delay in the “context” prevents a direct application of the elliptical potential lemma (carpentier2020elliptical), a staple in bandit regret analyses. We resolve this by adapting the analysis to our timing, recovering an appropriate potential-type bound (see Lemma C.2) that enables regret control despite the delayed feedback structure, which arises as a natural consequence of the competitive multi-seller set-up we consider in this paper.
@@ 1.2 Contributions
We summarize our main contributions and position them relative to prior works.
General monotone single–index demand. We model each seller’s mean demand as $\lambda_{i}(\mathbf{p})=\mu_{i}(\langle\boldsymbol{\theta}_{i,0},\mathbf{p}\rangle)$, where $\mu_{i}$ is a known and increasing link function but $\boldsymbol{\theta}_{i,0}$ is unknown. This framework strictly generalizes the linear model of li2024lego, capturing a broader range of nonlinear demand patterns. In contrast to li2024lego; bracale2025revenue; goyal2023learning, our analysis applies seamlessly to both binary (purchase/no-purchase) and real-valued (continuous) demand outcomes, thereby unifying common revenue models within a single learning procedure.
No forced i.i.d. exploration phase. Prior work relies on a coordinated, front-loaded exploration phase across all sellers (li2024lego; bracale2025revenue), in which i.i.d. randomized prices are posted to facilitate estimation. While effective in theory, the length of these “artificial” exploration phases depends on unknown constants, making them very difficult to calibrate in practice. Our policy removes the need for such an explicit exploration phase: sellers continuously learn while pricing, yet still achieve optimal statistical guarantees. This makes our method more practical and better aligned with realistic market conditions.
No-regret algorithm. We establish an upper bound on the total expected individual regret of order $O(N^{2}\sqrt{T}\log(T))$ per seller, up to additive constants. Moreover, we show that the price vector $\mathbf{p}^{(t)}$ satisfies the following bound $\sum_{t=1}^{T}\mathbb{E}[\|\mathbf{p}^{(t)}-\mathbf{p}^{*}\|_{2}^{2}]=O(N^{2}\sqrt{T}\log(T))$, where $\mathbf{p}^{*}$ is the NE. Thus, our analysis not only guarantees near–optimal regret, but also provides quantitative control over deviation of play from equilibrium.
@@ 2 Problem Setting
For all $t\in[T]$, the $i$-th seller’s price is denoted as $p_{i}^{(t)}\in\mathcal{P}_{i}\triangleq[\underline{p_{i}},\bar{p}_{i}]$, with price bounds $\underline{p_{i}}<\bar{p}_{i}$ and $\underline{p_{i}},\bar{p}_{i}\in[0,+\infty)$. Every seller knows their own price domain $\mathcal{P}_{i}$ but not the competing ones $\{\mathcal{P}_{j}\}_{j\in[N]\setminus\{i\}}$. However, only $B_{p}=(\sum_{i\in[N]}\bar{p}_{i}^{2})^{1/2}$ is known to all the sellers.
Let $\mathbf{p}_{-i}^{(t)}\triangleq(p_{j}^{(t)})_{j\in[N]\setminus\{i\}}$ denote the $i$-th competitors’ prices at time $t$ and $\mathbf{p}^{(t)}\triangleq(p_{j}^{(t)})_{j\in[N]}$ denote the joint prices. Only for the purpose of the analysis we define $\mathcal{P}\triangleq\prod_{i\in[N]}[\underline{p_{i}},\bar{p}_{i}]$ which denotes the support of joint prices and $\mathcal{P}_{-i}=\prod_{j\in[N]\setminus\{i\}}\mathcal{P}_{j}$ the support of the competitors’ prices. The general dynamic is as follows: at time $t=0$, each seller $i$ selects any initial price $p_{i}^{(0)}\in\mathcal{P}_{i}$ and everyone observes their own realized demand $y_{i}^{(0)}$ according to (2). The price vector $\mathbf{p}^{(0)}$ is made public. Let $\mathcal{H}^{(0)}_{i}=\{(\mathbf{p}^{(0)},y_{i}^{(0)})\}$ be the data collected by seller $i$. For each sales round $t=1,\ldots,T$:
@@
    Each seller $i$ sets their price $p_{i}^{(t)}$ based on their data history $\mathcal{H}_{i}^{(t-1)}$.
@@
    Each seller $i$ observes their (empirical) revenue $r_{i}=p_{i}^{(t)}y_{i}^{(t)}$.
@@
We consider a generalization of the monopolistic binary response model considered by fan2024policy; bracale2025dynamic (and corresponding references). Each customer arrives at time $t$ and, given the price vector posted by sellers $\mathbf{p}^{(t)}$, samples a valuation for each seller $i$
@@
Here we describe the model by li2024lego. The observed demand at time $t$ for seller $i$ is assumed to have the form
@@
We now present the assumptions on $\mu_{i}$ and the parameters $\boldsymbol{\theta}_{i,0}$. We assume that the average demand $\lambda_{i}$ is non-negative and $\partial_{p_{i}}\lambda_{i}<0$ among all values of $\{\boldsymbol{\theta}_{i,0}\}_{i\in[N]}$; a similar assumption is found in birge2024interfere; li2024lego; bracale2025revenue. The above conditions hold if $\mu_{i},\mu_{i}^{\prime}>0$ and $\beta_{i,0}>0$, which are explicitly stated in 2.4 and subsequent 2.5.
@@
Requiring $\mu_{i}^{\prime}$ to be bounded away from zero is standard: it prevents degeneracy in the curvature of the log-likelihood (equivalently, keeps the Fisher information well-conditioned) and underpins the concentration inequalities used in our analysis; similar assumptions can be found in russac2019weighted; russac2020algorithms; balabdaoui2019least; groeneboom2018current. Let $B_{i}^{\prime\prime}>0$ be such that $|\mu_{i}^{\prime\prime}(u)|\leq B_{i}^{\prime\prime}$ for all $u\in\mathcal{U}_{i}$, which exists because $\mu_{i}\in\mathrm{C}^{2}(\mathcal{U}_{i})$.
@@
Before proceeding to the next section, we highlight a key distinction between our framework and the MARL models discussed in Section 1.2. In most MARL settings, the realized rewards of all agents – in our setting the realized demands of competing sellers – are observed by every participant. By contrast, our model assumes that these realized demands remain private. This privacy has an important implication: no seller can infer the policy of their competitors. To see this, recall that the demand model for seller $j$ is $\mathbb{E}[y_{j}^{(t)}|\mathbf{p}^{(t)}]=\mu_{j}(\langle\boldsymbol{\theta}_{j,0},\mathbf{p}^{(t)}\rangle)$ where $\boldsymbol{\theta}_{j,0}$ and $\mu_{j}$ are the parameter vector and link function of seller $j$, respectively, both unknown to competitors. At each time $t$, a competitor seller $i\neq j$ only observes $\mathcal{H}_{i}^{(t)}=\cup_{s\leq t}\{(y_{i}^{(s)},\mathbf{p}^{(s)})\}$ but does not have access to the demand realizations $\{y_{j}^{(s)}\}_{s\leq t}$ of seller $j$. Thus, seller $i$ has no information from which to infer their competitor parameters $(\boldsymbol{\theta}_{j,0},\mu_{j})$, and hence predict their pricing policy.

### v4

# Supplementary for Online Price Competition under Generalized Linear Demands

@@ Abstract
We study sequential price competition among $N$ sellers, each influenced by the pricing decisions of their rivals. Specifically, the demand function for each seller $i$ follows the single index model $\lambda_{i}(\mathbf{p})=\mu_{i}(\langle\boldsymbol{\theta}_{i,0},\mathbf{p}\rangle)$, with known increasing link $\mu_{i}$ and unknown parameter $\boldsymbol{\theta}_{i,0}$, where the vector $\mathbf{p}$ denotes the vector of prices offered by all the sellers simultaneously at a given instant. Each seller observes only their own realized demand – unobservable to competitors – and the prices set by rivals. Our framework generalizes existing approaches that focus solely on linear demand models. We propose a novel decentralized policy, PML-GLUCB, that combines penalized MLE with an upper-confidence pricing rule, removing the need for front-loaded exploration phases across sellers — which is integral to previous linear models — and accommodating both binary and real-valued demand observations. Relative to a dynamic benchmark policy, each seller achieves $O(N^{2}\sqrt{T}\log(T))$ regret, which essentially matches the optimal rate known in the linear setting. A significant technical contribution of our work is the development of a variant of the elliptical potential lemma — typically applied in single-agent systems — adapted to our competitive multi-agent environment.
Machine Learning, ICML
@@ 1 Introduction
A pricing strategy aims to determine prices that maximize current profits. Much of the learning-to-price literature, however, has traditionally focused on monopolistic settings (Fan et al., 2024; Javanmard and Nazerzadeh, 2019; Golrezaei et al., 2019; Bracale et al., 2025b), effectively treating demand as independent of competitors’ actions. In many real-world markets, this assumption is unrealistic. Markets typically feature multiple sellers offering products that differ in quality and price, each seeking to maximize its own revenue. In such environments, a seller’s profit is inherently affected by the pricing decisions of its competitors. As a result, firms must continuously revise their prices to anticipate and respond to rivals’ moves. Motivated by these considerations, recent work has moved towards the study of pricing policies in multi-seller environments, explicitly modeling strategic interactions among competing firms (Li et al., 2024; Bracale et al., 2025a; Yang et al., 2024; Goyal et al., 2023; Zhu et al., 2025). In such settings, each seller seeks to maximize its own revenue, which inherently depends on the pricing policies adopted by its competitors. This interdependence naturally leads to the concept of a Nash equilibrium: a profile of prices at which no seller can improve its payoff through a unilateral deviation, given the strategies of the others. Building on this framework, existing works (Li et al., 2024; Bracale et al., 2025a) propose sequential pricing mechanisms designed to learn equilibrium behavior over time, and establish sublinear regret guarantees relative to the equilibrium benchmark [^1] In this paper, we use the terms sublinear regret and no regret interchangeably to indicate that the cumulative regret grows sub-linearly with time, that is, $\mathrm{Reg}(T)=o(T)$ as $T\to\infty$. and converge to a Nash equilibrium (NE) (Li et al., 2024; Bracale et al., 2025a). Yet two practical issues remain:
  1. 1.
    Existing non-monopolistic approaches (Li et al., 2024; Bracale et al., 2025a) typically decouple exploration” from optimization” and rely on an explicit, front-loaded exploration phase with an i.i.d. pricing policy. However, as noted by Jin et al. (2021), classical results such as Garivier et al. (2016) show that the explore-then-commit strategy can be asymptotically suboptimal as the horizon grows, and can therefore be dominated by fully sequential policies such as Upper Confidence Bound (UCB). Moreover, in some dynamic-pricing implementations, the required length of the exploration phase depends on unknown problem-dependent constants (Bracale et al., 2025a), which complicates deployment in practice.
  2. 2.
    The demand is often forced to be linear (or narrowly parameterized) (Bertsimas and Perakis, 2006; Guda and Subramanian, 2026; Li et al., 2024; Goyal et al., 2023), limiting the ability to capture nonlinear demand patterns commonly observed in real markets (Gallego et al., 2006; Wan et al., 2022, 2023).
To address these limitations, we adopt a generalized linear (single-index) demand framework, which accommodates a broader range of nonlinear demand responses and naturally covers both binary and continuous demand outcomes, and propose a pricing policy that operates without a dedicated initial i.i.d. pricing-design exploration phase. We prove that each seller’s regret remains on the order of $\widetilde{O}(\sqrt{T})$ against a dynamic benchmark, where $\widetilde{O}$ suppresses logarithmic factors.
@@ 1.1 Related Literature
Over the past two decades, there has been substantial progress in demand learning under competition. Early work by Kirman (1975) analyzed symmetric two-seller competition with linear demand. Asymmetry and noise were incorporated by Cooper et al. (2015), who let each seller learn independently (effectively as a monopolist) and established equilibrium under known price sensitivities. In contrast, Li et al. (2024) studied dynamic pricing with unknown price sensitivities under a linear demand model and achieved the optimal $\widetilde{O}(\sqrt{T})$ regret rate. In their setup, seller $i$’s expected demand is
@@
where $N]=\{1,2,\dots,N\}$ indexes the sellers and the parameters $(\alpha_{i,0},\beta_{i,0},\gamma_{ij,0})$ are unknown with $\beta_{i,0}>0$. The condition $\beta_{i,0}>0$ captures the standard economic assumption that the demand of seller $i$ decreases as their own price $p_{i}$ increases. However, nonlinear demand models are often more realistic and have been explored in several applications (Gallego et al., 2006; Wan et al., 2022, 2023). For instance, Gallego et al. (2006) considered the mean demand function $\lambda_{i}(\mathbf{p})=a_{i}\left(p_{i}\right)/(\sum_{j}a_{j}\left(p_{j}\right)+\kappa)$, for some known increasing functions $a_{i}$ and real value $\kappa\in(0,1]$. They establish the existence of a NE, but they do not propose an online learning procedure.
In contrast to prior studies (Guda and Subramanian, 2026; Li et al., 2024; Kachani et al., 2007; Gallego et al., 2006), we analyze a more general monotone single-index expected-demand model:
@@
for each $i\in[N]$, where $\mathbf{p}_{-i}=(p_{j})_{j\in[N]\setminus\{i\}}$, $\mathbf{p}=(p_{j})_{j\in[N]}$ and $\boldsymbol{\theta}_{i,0}$ is a vector of dimension $N$ with $i$-th entry equals to $\beta_{i,0}$ (where $\beta_{i,0}>0$) and the rest of the entries are the values of $\boldsymbol{\gamma}_{i,0}$ ordered. The link function $\mu_{i}$ is assumed to be increasing, capturing the same economic intuition that demand decreases with the seller’s own price $p_{i}$. Here $\mu_{i}$ is known to seller $i$ (but unknown to competitors) and the parameter $\boldsymbol{\theta}_{i,0}$ is unknown (to everyone). This model generalizes the linear demand model by Li et al. (2024); specifically, $\mu_{i}(u)=u+\alpha_{i,0}$ recovers their formulation.
@@
A key modeling assumption in MARL algorithms is full or partial observability of other agents’ rewards or payoffs (Hu and Wellman, 2003; Zhang et al., 2021; Yang and Wang, 2020). For instance, the canonical formulation of Hu and Wellman (2003) assumes that each agent can observe the immediate reward of all players and use this to update joint policy. Similarly, Zhang et al. (2021) and Yang and Wang (2020) rely on either joint reward observability or access to sufficient statistics (e.g., payoff signals or utility vectors) that allow agents to infer the equilibrium structure. This assumption is critical: observing other agents’ realized rewards enables an agent to estimate the full payoff landscape and to learn best-response dynamics with respect to the joint reward function.
Our setting fundamentally differs from this MARL paradigm in both information structure and observability. In the sequential pricing game we study, each seller observes only their own reward (or demand value) and rivals’ posted prices – never competitors’ demands or revenues. Because demand functions are private and unobservable, a seller cannot recover rivals’ payoff functions or utility gradients, so classical MARL techniques are not directly applicable. Nevertheless, we design a policy whose dynamics need not converge to a Nash equilibrium, yet whose individual regret satisfies $\widetilde{O}(\sqrt{T})$. This makes our approach better aligned with real-world competitive pricing, where revenue data is proprietary, and competitors’ demand is hidden.
@@
One may view $\mathbf{p}_{-i}^{(t)}$ at any instant $t$ as the “context” for seller $i$, linking our problem to contextual bandits (Chu et al., 2011; Agrawal and Goyal, 2013; Chakraborty et al., 2023; Kim et al., 2023) and to the nonlinear Cournot structure (Cournot, 1838; Saloner, 1987). That said, our model departs from the standard contextual bandit framework in two key ways. First, in the bandit literature, the context is typically exogenous — sampled independently of the agent’s actions — whereas here the “context” $\mathbf{p}_{-i}^{(t)}$ evolves as a function of past prices (and thus past demand noise realizations). Second, classical contextual bandits assume the context is observed before the action is chosen; in our market, sellers post prices simultaneously, so seller $i$ observes $\mathbf{p}_{-i}^{(t)}$ only after selecting $p_{i}^{(t)}$. This delayed observation prevents a direct application of standard contextual bandit analyses, in particular arguments based on the elliptical potential lemma (Carpentier et al., 2020), which is central to many regret proofs. We address this timing mismatch by adapting the analysis to our simultaneous-move setting, deriving an appropriate elliptical-type bound (see Lemma C.2) that yields regret control despite the delayed “context”.
@@ 1.2 Contributions
We summarize our main contributions.
[…]
(2) Generalize linear demand with private feedback. We study a competitive demand-learning model of the form $\lambda_{i}(\mathbf{p})=\mu_{i}(\langle\boldsymbol{\theta}_{i,0},\mathbf{p}\rangle),$ where $\mu_{i}$ is a known increasing link and $\boldsymbol{\theta}_{i,0}$ is unknown. This strictly generalizes the linear models (Li et al., 2024; Guda and Subramanian, 2026) and, in contrast to prior works, naturally covers both (i) binary outcomes (purchase / no-purchase) (Goyal et al., 2023; Golrezaei et al., 2019; Javanmard, 2017; Javanmard and Nazerzadeh, 2019; Fan et al., 2024; Bracale et al., 2025b) and (ii) continuous demand observations (Bertsimas and Perakis, 2006; Li et al., 2024; Bracale et al., 2025a) within the same analysis. Crucially, our information structure matches real markets: each seller observes only their own realized demand/revenue and competitors’ posted prices, not competitors’ outcomes, ruling out standard multi-agent RL assumptions (as noted above in Section 1.1) and requiring new proof techniques.
[…]
@@ 2 Problem Setting
For all $t\in[T]$, the $i$-th seller’s price is denoted as $p_{i}^{(t)}\in\mathcal{P}_{i}\triangleq[\underline{p_{i}},\bar{p}_{i}]$, with price bounds $0<\underline{p_{i}}<\bar{p}_{i}<+\infty$. Every seller knows their own price domain $\mathcal{P}_{i}$ but not the competing ones $\{\mathcal{P}_{j}\}_{j\in[N]\setminus\{i\}}$. However, only $B_{p}=(\sum_{i\in[N]}\bar{p}_{i}^{2})^{1/2}$ is known to all the sellers.
Let $\mathbf{p}_{-i}^{(t)}\triangleq(p_{j}^{(t)})_{j\in[N]\setminus\{i\}}$ denote the $i$-th seller’s competitors’ prices at time $t$ and $\mathbf{p}^{(t)}\triangleq(p_{j}^{(t)})_{j\in[N]}$ denote the joint prices. Only for the purpose of the analysis we define $\mathcal{P}\triangleq\prod_{i\in[N]}[\underline{p_{i}},\bar{p}_{i}]$ which denotes the support of joint prices and $\mathcal{P}_{-i}=\prod_{j\in[N]\setminus\{i\}}\mathcal{P}_{j}$ the support of the competitors’ prices. The general dynamic is as follows: at time $t=0$, each seller $i$ selects any initial price $p_{i}^{(0)}\in\mathcal{P}_{i}$ and everyone observes their own realized demand $y_{i}^{(0)}$ according to (2). The price vector $\mathbf{p}^{(0)}$ is made public. Let $\mathcal{H}^{(0)}_{i}=\{(\mathbf{p}^{(0)},y_{i}^{(0)})\}$ be the data collected by seller $i$. For each sales round $t=1,\ldots,T$:
@@
    Each seller $i$ sets $p_{i}^{(t)}$ based on their history $\mathcal{H}_{i}^{(t-1)}$.
@@
    Each seller $i$ observes their revenue $r_{i}=p_{i}^{(t)}y_{i}^{(t)}$.
@@
We consider a generalization of the monopolistic binary response model considered by Fan et al. (2024); Bracale et al. (2025b) (and corresponding references). Each customer arrives at time $t$ and, given the price vector posted by sellers $\mathbf{p}^{(t)}$, samples a valuation for each seller $i$
@@
Here we describe the model by Li et al. (2024). The observed demand at time $t$ for seller $i$ is assumed to have the form
@@
We now present the assumptions on $\mu_{i}$ and the parameters $\boldsymbol{\theta}_{i,0}$. We assume that the average demand $\lambda_{i}$ is non-negative and $\partial_{p_{i}}\lambda_{i}<0$ among all values of $\{\boldsymbol{\theta}_{i,0}\}_{i\in[N]}$; a similar assumption is found in Birge et al. (2024); Li et al. (2024); Bracale et al. (2025a). The above conditions hold if $\mu_{i},\mu_{i}^{\prime}>0$ and $\beta_{i,0}>0$, which are explicitly stated in 2.4 and subsequent 2.5.
@@
Requiring $\mu_{i}^{\prime}$ to be bounded away from zero is standard: it prevents degeneracy in the curvature of the log-likelihood (equivalently, keeps the Fisher information well-conditioned) and underpins the concentration inequalities used in our analysis; similar assumptions can be found in Russac et al. (2019, 2020); Balabdaoui et al. (2019); Groeneboom and Hendrickx (2018). Let $B_{i}^{\prime\prime}>0$ be such that $|\mu_{i}^{\prime\prime}(u)|\leq B_{i}^{\prime\prime}$ for all $u\in\mathcal{U}_{i}$, which exists because $\mu_{i}\in\mathrm{C}^{2}(\mathcal{U}_{i})$.
@@
Before proceeding to the next section, we highlight a key distinction between our framework and the MARL models discussed in Section 1.2. In most MARL settings, the realized rewards of all agents – in our setting, the realized demands of competing sellers – are observed by every participant. By contrast, our model assumes that these realized demands remain private. This privacy has an important implication: no seller can infer the policy of their competitors. To see this, recall that the demand model for seller $j$ is $\mathbb{E}[y_{j}^{(t)}|\mathbf{p}^{(t)}]=\mu_{j}(\langle\boldsymbol{\theta}_{j,0},\mathbf{p}^{(t)}\rangle)$ where $\boldsymbol{\theta}_{j,0}$ and $\mu_{j}$ are the parameter vector and link function of seller $j$, respectively, both unknown to competitors. At each time $t$, a competitor seller $i\neq j$ only observes $\mathcal{H}_{i}^{(t)}=\cup_{s\leq t}\{(y_{i}^{(s)},\mathbf{p}^{(s)})\}$ but does not have access to the demand realizations $\{y_{j}^{(s)}\}_{s\leq t}$ of seller $j$. Thus, seller $i$ has no information from which to infer their competitor parameters $(\boldsymbol{\theta}_{j,0},\mu_{j})$, and hence predict their pricing policy.